# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 02 Mar 2026 v1

@@ Abstract
We develop a Fisher-geometric theory of stochastic gradient descent (SGD) in which mini-batch noise is an intrinsic, loss-induced matrix—not an exogenous scalar variance. Under exchangeable sampling, the mini-batch gradient covariance is pinned down (to leading order) by the projected covariance of per-sample gradients: it equals projected Fisher information for well-specified likelihood losses and the projected Godambe (sandwich) matrix for general M-estimation losses. This identification forces a diffusion approximation with Fisher/Godambe-structured volatility (effective temperature $\tau=\eta/b$) and yields an Ornstein–Uhlenbeck linearization whose stationary covariance is given in closed form by a Fisher–Lyapunov equation. Building on this geometry, we prove matching minimax upper and lower bounds of order $\Theta(1/N)$ for Fisher/Godambe risk under a total oracle budget $N$; the lower bound holds under a martingale oracle condition (bounded predictable quadratic variation), strictly subsuming i.i.d. and exchangeable sampling. These results imply oracle-complexity guarantees for $\varepsilon$-stationarity in the Fisher dual norm that depend on an intrinsic effective dimension and a Fisher/Godambe condition number rather than ambient dimension or Euclidean conditioning. Experiments confirm the Lyapunov predictions and show that scalar temperature matching cannot reproduce directional noise structure.
@@ 1 Introduction
Stochastic gradient descent (SGD) is the computational workhorse behind a large class of operations research and management science problems, including simulation optimization, large-scale stochastic programming, and data-driven estimation in service and supply-chain systems. In these settings, sampling effort is a decision variable: each iterate trades off additional replications (or scenarios) against the number of updates that can be executed under a fixed budget. Several widely observed phenomena depend on this tradeoff: small mini-batches often dominate large batches in wall-clock efficiency; constant-step SGD exhibits steady-state “error floors” and implicit regularization; and plain SGD frequently behaves as if it were curvature-aware despite using only first-order information.
This paper offers a unified account of these phenomena by identifying a structural property of the stochastic-gradient oracle that is often treated as exogenous in stochastic approximation: mini-batch gradient noise has a loss-induced matrix geometry. Under exchangeable sampling, the leading-order covariance of the mini-batch average is determined by the population covariance of per-sample gradients projected onto identifiable directions. In likelihood settings this matrix is Fisher information; for general losses it is the corresponding Godambe/sandwich geometry. Once this intrinsic matrix is identified, (i) the diffusion approximation used to analyze constant-step SGD becomes structurally determined, (ii) the natural metric for rates is Fisher/Godambe rather than Euclidean, and (iii) batch size admits principled operations-research (OR) design rules. Our central result is thus an identification, not merely a bound: the noise covariance is pinned down by the sampling mechanism, and this single structural fact forces the diffusion coefficient, determines the natural convergence metric, and yields tight oracle-complexity guarantees.
While diffusion analyses of constant-step SGD are classical, most treatments either assume isotropic gradient noise or summarize noise effects through scalar quantities such as trace-level variance or Euclidean error. In contrast, we identify the full matrix geometry of mini-batch noise under exchangeable sampling and show that this structure alone determines the diffusion coefficient, the natural convergence metric, and the resulting oracle complexity. The contribution is therefore structural: the geometry is pinned down by the sampling mechanism itself.
Formally, the paper establishes the following results:
  1. 1.
    Identification of intrinsic mini-batch noise geometry. Under exchangeable sampling, the leading-order mini-batch gradient covariance is uniquely determined by the projected covariance of per-sample gradients. In likelihood models this reduces to Fisher information; in general M-estimation it becomes the Godambe matrix. Thus the diffusion coefficient is not a modeling assumption but a structural consequence of the sampling mechanism.
  2. 2.
    Fisher-structured stationary law. The induced diffusion admits an Ornstein–Uhlenbeck linearization whose stationary covariance solves a Fisher–Lyapunov equation, yielding closed-form equilibrium structure and explicit directional variance formulas.
  3. 3.
    Minimax-optimal Fisher/Godambe risk bounds. Under a martingale oracle condition (bounded predictable quadratic variation), we prove matching upper and lower bounds of order $\Theta(1/N)$ for Fisher/Godambe risk under a total oracle budget $N$.
  4. 4.
    Oracle complexity in intrinsic dimension. The resulting complexity guarantees for $\varepsilon$-stationarity in the Fisher dual norm depend on an intrinsic effective dimension and a Fisher/Godambe condition number rather than ambient dimension or Euclidean conditioning.
@@ 1.1 Motivation: Simulation Optimization as Sampling Control
followed by $\theta^{+}=\theta-\eta\,g_{B}(\theta)$. Classical variance-reduction intuition focuses on the scalar variance of a component or of a scalarized performance measure, suggesting that larger $b$ should always help. In practice, under fixed compute or sample budgets, small $b$ often wins.
The key reason is geometric: the noise in $g_{B}(\theta)$ is not scalar. Its covariance is a matrix that encodes which directions are intrinsically informative and which are weakly identified by the underlying experiment. Our results show that, to leading order, this covariance scales as $b^{-1}G^{*}(\theta)$, where $G^{*}(\theta)$ is the projected covariance of per-sample gradients. For likelihood-based objectives, $G^{*}(\theta)$ reduces to projected Fisher information; for general M-estimation and proper-scoring objectives it is the corresponding Godambe-aligned object. This provides the statistical foundation for interpreting SGD as a controlled diffusion whose “temperature” is $\tau=\eta/b$.
OR problem classes where the theory applies.
The sampling-control perspective above is not limited to simulation optimization. Three canonical OR problem classes share the same structure: (i) simulation optimization (queuing, inventory, revenue management), where each mini-batch gradient requires $b$ simulation replications and the replication budget is the binding constraint; (ii) stochastic programming / SAA, where scenario-based gradient estimates are drawn from an empirical distribution and the number of scenarios per iterate is a design variable subject to a total scenario budget; and (iii) online M-estimation in service systems, where per-transaction loss gradients arrive in a stream and the batch window trades off estimation quality against decision latency. In each case, sampling effort is explicit, $b$ is small relative to $d$, and the question “how many samples per iterate?” is operationally binding. The theory developed below applies uniformly to all three classes.
@@ 1.2 Core Insights
Before turning to formal statements, we summarize the main idea informally. When SGD computes a mini-batch gradient by averaging $b$ per-sample gradients, the resulting noise is not “generic randomness.” Its covariance—the matrix describing which directions are noisy and which are quiet—is determined by the loss function and the data-generating process. Specifically, directions in parameter space where the data carry more information (in the sense of Fisher information) naturally produce larger gradient fluctuations.
This has three immediate consequences that do not require any SDE machinery to appreciate:
@@
    The noise is structured, not spherical. If one doubles the batch size, noise is reduced equally in all directions, but the shape of the noise ellipsoid is fixed by the statistical problem. This shape is the Fisher information matrix (for likelihood losses) or its generalization, the Godambe matrix (for arbitrary losses).
@@
    The right metric for measuring SGD error is statistical, not Euclidean. Since the noise is Fisher-shaped, measuring convergence in the Fisher metric (which weights directions by their statistical informativeness) gives tighter and more interpretable bounds than the Euclidean metric.
@@
    Batch size is a design variable, not just a hyperparameter. Because batch size controls the “temperature” $\tau=\eta/b$ of a diffusion process, it directly governs the Fisher/Godambe-weighted stationary covariance through the Lyapunov balance. This makes $b$ a first-class design lever: under a fixed sampling budget $N=Tb$, the oracle complexity (Theorem 7.3) scales with the intrinsic dimension $d_{\operatorname{eff}}$ and condition number $\kappa_{F}$, clarifying when small batches are near-optimal rather than merely convenient.
@@ 1.3 What We Show
(i) Alignment from sampling theory.
@@
where $G^{*}(\theta)$ is the projected gradient-noise covariance (Section 4). In particular, for correctly specified likelihood losses $G^{*}(\theta)=F^{\star}(\theta)$ (Fisher alignment). That is, the noise in a mini-batch gradient is not arbitrary—to leading order, it is proportional to the Fisher/Godambe matrix, which encodes how informative the data are about the parameters.
@@
In the small-stepsize regime with $\tau=\eta/b$ held fixed, the SGD recursion admits a diffusion approximation whose diffusion covariance is $\tau\,G^{*}(\theta)$ (Section 5). Near a nondegenerate critical point, the limiting SDE linearizes to an OU process whose stationary covariance $\Sigma_{\infty}$ solves the Lyapunov equation
@@
making explicit that steady-state fluctuations are governed by the ratio of curvature to Fisher/Godambe noise.
(iii) Matching Fisher-metric rates and minimax lower bounds.
We analyze SGD performance in the frozen Fisher/Godambe geometry at $\theta^{\star}$. We prove $1/T$ upper bounds for mean-square error in the Fisher metric and match them with information-theoretic lower bounds (via van Trees), showing the rate and its dependence on the intrinsic geometry are minimax tight (Sections 6 and 7).
(iv) Oracle complexity in statistical rather than Euclidean conditioning.
Our oracle complexity bounds scale with a Fisher/Godambe condition number and an effective dimension, isolating statistical difficulty from purely Euclidean conditioning. This clarifies when classical Hessian-based condition numbers can be misleading for statistical objectives.
@@
The framework above connects batch-size design to Fisher/Godambe geometry; Section 9 discusses extensions to misspecification, dependence, and operational effort allocation.
@@
Theorem 4.3 — Mini-batch covariance $\tfrac{1}{b}\,G^{\star}(\theta)$ (Fisher under correct specification). Theorem 5.4 — SGD $\Rightarrow$ SDE with Fisher/Godambe-structured volatility at temperature $\tau\!=\!\eta/b$. Thm. 5.7, Cor. 5.8 — OU linearization at $\theta^{\star}$; stationary covariance solves a Lyapunov equation. Thms. 6.4, 6.8 — Matching $1/(Tb)$ upper and lower bounds in the Fisher metric. Theorem 7.3 — Oracle complexity $N\!=\!\widetilde{\Theta}(\kappa_{F}\,d_{\mathrm{eff}}/\varepsilon^{2})$ for Fisher-dual stationarity.
1.4 Positioning and Organization
Our analysis connects classical stochastic approximation (Robbins and Monro, 1951; Kushner and Yin, 2003) with diffusion views of SGD (Mandt et al., 2017; Li et al., 2017) and information geometry (Amari, 1998), while emphasizing the operations-research-relevant regime where sampling effort is explicit and often $b\ll d$. Unlike diffusion analyses that posit a fixed gradient-noise covariance near stationarity, we derive the covariance structure directly from exchangeable sampling. In our framework, the Fisher/Godambe geometry is not an auxiliary modeling assumption but a consequence of the sampling mechanism itself.
Section 2 reviews related work and positions the contribution. Section 3 defines the oracle model, the geometric objects ($G^{*}$, Fisher, Godambe), and the local regularity conditions. Section 4 derives the mini-batch covariance alignment from sampling theory and exchangeability. Section 5 builds the diffusion approximation and local OU/Lyapunov structure on this alignment. Section 6 proves matching upper and lower bounds in Fisher geometry, and Section 7 states oracle complexity in the Fisher/Godambe metric. Section 8 provides numerical validation via two key diagnostics: (i) the Lyapunov/OU stationary risk prediction is accurate—the empirical Fisher-metric risk plateaus at the predicted $\mathrm{Tr}(G^{\star}H^{-1})$ level; and (ii) scalar temperature matching fails directionally—a rotation sweep reveals cross-covariance structure that no isotropic surrogate can reproduce. Section 9 discusses extensions to misspecification, dependence, and OR-specific design rules.
Reading guide.
Readers primarily interested in practical implications may start with Section 7 (oracle complexity) and Section 8 (numerical validation), referring back to Sections 45 for the underlying alignment and diffusion theory as needed. Readers interested in the theoretical contributions should follow the paper linearly. The main text includes proofs of all spine results; auxiliary lemmas and concentration tools are collected in the appendices.
@@ 2 Literature Review and Positioning
This paper sits at the intersection of stochastic approximation, diffusion approximations for SGD, information geometry, and oracle-complexity theory, with a deliberate focus on operations research settings where sampling effort is an explicit decision variable. In these settings, the mini-batch size $b$ is a budget-constrained design choice: our results convert it into an effective temperature control via the ratio $\tau=\eta/b$, providing actionable design rules for practitioners allocating simulation effort. The unifying thread is that SGD is a statistical experiment: the stochastic gradient oracle is generated by sampling, and its noise therefore has an intrinsic geometry determined by the loss and the data model.
@@
To make the contribution transparent, we organize prior work around two recurring degrees of freedom: (i) whether the mini-batch noise covariance is assumed (exogenous) or derived from the sampling mechanism; and (ii) whether performance and rates are stated in a Euclidean metric or in a statistically meaningful metric (Fisher/Godambe). Table 1 summarizes the resulting taxonomy of primitives and contrasts, and provides pointers to the corresponding sections and theorems in this paper.
2.1 Stochastic Approximation: Rates with Exogenous Noise Geometry
Classical stochastic approximation studies recursions of the form
$\theta_{t+1}=\theta_{t}-\eta_{t}\{h(\theta_{t})+\xi_{t+1}\},$ (5)
where $h$ is a vector field and $\xi_{t}$ is a martingale-difference noise process. Under stability and integrability conditions, $\theta_{t}\to\theta^{\star}$ with $h(\theta^{\star})=0$, and with additional smoothness one obtains a $\sqrt{t}$-CLT whose covariance is characterized by a Lyapunov equation involving the Jacobian $J(\theta^{\star})$ and a long-run noise covariance $\Gamma$ (Robbins and Monro, 1951; Dvoretzky, 1956; Kushner and Yin, 2003). Optimal finite-time rates under averaging were established by Polyak and Juditsky (1992).
This framework is foundational, but for stochastic gradients it leaves the key object $\Gamma$ exogenous. In SGD, the noise is endogenous: it is created by mini-batching and therefore inherits structure from the sampling design and the statistical objective. Our first contribution is to make this structure explicit. As summarized in the first row of Table 1, Theorem 4.3 derives the leading-order covariance of the mini-batch mean from exchangeable sampling (with finite-population correction), identifying the intrinsic matrix $G^{\star}(\theta)$ (Fisher for likelihood losses, Godambe/sandwich more generally) that governs subsequent diffusion and rate results.
2.2 Diffusion Approximations: Which Diffusion Matrix?
A large literature interprets constant-step SGD as an approximate discretization of an SDE and uses its stationary distribution to explain implicit regularization and generalization behavior (Mandt et al., 2017; Li et al., 2017; Smith et al., 2021). A common template is
$d\theta_{t}=-\nabla L(\theta_{t})\,dt+\sqrt{2\tau}\,\mathbf{C}(\theta_{t})^{1/2}dW_{t},\qquad\tau=\eta/b,$ (6)
with multiple proposed specifications for $\mathbf{C}(\theta)$ (isotropic, Hessian-based, Fisher-based, or empirically estimated).
Our results eliminate this modeling degree of freedom in the exchangeable/i.i.d. baseline. As summarized in Table 1, the diffusion matrix is not a free choice: it is implied by the mini-batch design and equals (to leading order) the projected gradient-noise covariance $G^{\star}(\theta)$; see Theorem 4.3. This pins down the diffusion approximation in Theorem 5.4 and yields a local OU regime with a Lyapunov balance governed by $G^{\star}$ and temperature $\tau=\eta/b$; see Corollary 5.8.
Contrast with existing diffusion analyses.
Mandt et al. (2017) treat the diffusion coefficient as a tunable modeling choice calibrated post hoc; Li et al. (2017) derive modified equations that capture discretization bias but do not identify the covariance with a population-level statistical object; Smith et al. (2021) connect implicit regularization to noise structure but work in a Hessian-scaling regime rather than deriving the matrix from sampling theory. Our contribution is complementary and more primitive: we show that, under standard mini-batch sampling, the diffusion matrix is forced to be $G^{\star}(\theta)$—a quantity determined entirely by the loss and the data model—before any discretization correction or regularization argument is applied. This identification is what enables the subsequent rate-matching bounds (Theorems 6.46.8) and the oracle-complexity guarantee (Theorem 7.3) in a statistically natural metric.
2.3 Scaling Regimes and the Small-Batch Operating Region
Recent rigorous diffusion-limit work studies joint regimes in which dimension, batch size, and step size scale together (Paquette et al., 2024). These results provide clean functional CLTs, but typically emphasize settings where $b$ grows proportionally with $d$.
Our analysis is complementary and closer to operations-research operating constraints. We keep $d$ fixed and treat $b$ as an explicit design variable, with particular attention to the common regime $b\ll d$. In this regime, it is not enough to track a scalar variance level; the matrix geometry of $G^{\star}(\theta)$ controls both transient behavior and risk. This viewpoint feeds directly into Fisher-metric finite-time rates (Section 6) and oracle complexity (Section 7; Theorem 7.3); see Table 1.
2.4 Information Geometry and Natural Gradients
Information geometry treats Fisher information as a Riemannian metric and motivates natural-gradient updates (Amari, 1998; Martens, 2020):
$\theta_{t+1}=\theta_{t}-\eta\,F(\theta_{t})^{-1}\nabla L(\theta_{t}).$ (7)
Natural gradients enjoy invariance and improved conditioning, but they require estimating and (approximately) inverting $F(\theta)$.
Our contribution is orthogonal: we identify where intrinsic statistical geometry enters plain SGD. The geometry appears implicitly through the diffusion term rather than the drift: mini-batch noise injects anisotropic fluctuations aligned with $G^{\star}(\theta)$, and the local OU equilibrium shows that stationary spread is governed by a Lyapunov balance rooted in the intrinsic statistical geometry (Corollary 5.8). This sharpens the empirical observation that SGD can behave as if it were “curvature-aware” without explicit preconditioning: SGD is not a natural-gradient method, but its noise is naturally shaped.
2.5 Minimax Rates and Information-Theoretic Limits
Lower bounds for stochastic first-order optimization are commonly stated in Euclidean norms under smoothness and strong convexity (Nemirovski et al., 2009; Agarwal et al., 2012). For statistical objectives, such bounds can be loose because Euclidean distance is not the right notion of local distinguishability.
We therefore state both upper and lower bounds in the Fisher/Godambe metric. Using a van Trees (Bayesian Cramér–Rao) argument, we derive a minimax lower bound in Fisher geometry (Theorem 6.8) and match it with a Fisher-Lyapunov upper bound (Theorem 6.4). This identifies the intrinsic drivers of complexity as statistical curvature and dimensionality in the induced metric (e.g., $\lambda_{\min}(F^{\star})$ and $d_{\operatorname{eff}}$), as summarized in Table 1.
2.6 Operations Research Context: Sampling Budgets and Effort Allocation
In simulation optimization, stochastic programming, and sample-average approximation (SAA), gradients are estimated from sampled scenarios and sampling cost is explicit (Fu, 2006; Shapiro et al., 2014; Chernoff, 1959). Batch size is therefore not merely a tuning knob: it is a decision variable subject to a budget.
Our framework connects this OR viewpoint to Fisher/Godambe geometry. Because batching controls temperature $\tau=\eta/b$, it directly shapes the diffusion amplitude and hence terminal risk in the appropriate metric. Section 9 outlines how this naturally leads to a covariance-control / effort allocation problem. Table 1 summarizes the positioning of our contributions.
2.7 Why Small Batches Can Work When $b\ll d$
The empirical preference for small batches persists even in very high-dimensional settings (Keskar et al., 2017; Shallue et al., 2019; McCandlish et al., 2018). Our results provide a principled explanation compatible with OR budgeting:
  1. 1.
    Throughput dominates precision under fixed sample budgets: $N$ total samples imply $T=N/b$ updates, so smaller $b$ buys more opportunities for contraction.
  2. 2.
    Geometry governs exploration: under Fisher/Godambe alignment, noise is anisotropic and concentrates along statistically flat directions, not merely “high variance” in Euclidean units.
  3. 3.
    Effective dimension can be much smaller than $d$: the relevant rates scale with $d_{\operatorname{eff}}$ rather than ambient dimension.
  4. 4.
    Operational constraints are not merely limitations: practical ceilings on $b$ can be near-optimal when viewed through information-efficiency per unit sampling cost.
The message is not that smaller batches are always better, but that the right batch size is a function of statistical geometry and budget, not only hardware or scalar variance reduction.
Across all of these threads, the contribution of this paper reduces to a single identification—the noise covariance of mini-batch SGD is determined by the loss and the sampling design—from which the diffusion limit, the rate-matching bounds, and the oracle complexity follow as necessary consequences.
@@ 2.7 Why Small Batches Can Work When $b\ll d$
Mini-batch noise covariance Exogenous / isotropic / heuristic $\Gamma$ Derived from sampling: $\mathrm{Cov}(g_{B})\propto b^{-1}G^{\star}(\theta)$; Fisher is a special case Sec. 4; Thm. 4.3
Diffusion matrix for SGD Postulated $\mathbf{C}(\theta)$ in SDEs Diffusion coefficient is forced by sampling theory: $G^{\star}(\theta)$ Sec. 5; Thm. 5.4
OU / Lyapunov equilibrium Heuristic OU with ad hoc covariance OU linearization with Lyapunov balance driven by $G^{\star}$ at temperature $\tau=\eta/b$ Sec. 5; Cor. 5.8
Metric for rates Euclidean distance; $\kappa_{H}$ dominates Fisher/Godambe metric is the natural risk Sec. 6; Thm. 6.4
Lower bounds Euclidean oracle lower bounds Information-theoretic lower bound in Fisher metric (van Trees) Sec. 6; Thm. 6.8

### 03 Apr 2026 v2

@@ Abstract
Classical stochastic-approximation analyses treat the covariance of stochastic gradients as an exogenous modeling input. We show that under exchangeable mini-batch sampling this covariance is identified by the sampling mechanism itself: to leading order it is the projected covariance of per-sample gradients. In well-specified likelihood problems this reduces locally to projected Fisher information; for general M-estimation losses the same object is the projected gradient covariance $G^{\star}(\theta)$, which together with the Hessian induces sandwich/Godambe geometry.
This identification—not the subsequent diffusion or Lyapunov machinery, which is classical once the noise matrix is given—is the paper’s main contribution. It endogenizes the diffusion coefficient (with effective temperature $\tau=\eta/b$), determines the stationary covariance via a Lyapunov equation whose inputs are now structurally fixed, and selects the identified statistical geometry as the natural metric for convergence analysis. We prove matching upper and lower bounds of order $\Theta(1/N)$ for risk in this metric under an oracle budget $N$; the lower bound is established first via a van Trees argument in the parametric Fisher setting and then extended to adaptive oracle transcripts under a predictable-information condition and mild conditional likelihood regularity. Translating these bounds into oracle complexity yields $\varepsilon$-stationarity guarantees in the Fisher dual norm that depend on an intrinsic effective dimension $d_{\mathrm{eff}}$ and a statistical condition number $\kappa_{F}$, rather than ambient dimension or Euclidean conditioning. Numerical experiments confirm the Lyapunov predictions at both continuous-time and discrete-time levels and show that scalar temperature matching cannot reproduce directional noise structure.
@@ 1 Introduction
Stochastic gradient descent (SGD) is the computational workhorse behind a large class of operations research and management science problems, including simulation optimization, large-scale stochastic programming, and data-driven estimation in service and supply-chain systems. In these settings, sampling effort is a decision variable: each iterate trades off additional replications (or scenarios) against the number of updates that can be executed under a fixed budget. Several widely observed phenomena depend on this trade-off: small mini-batches often dominate large batches in wall-clock efficiency; constant-step SGD exhibits steady-state “error floors” and implicit regularization; and plain SGD frequently behaves as if it were curvature-aware despite using only first-order information.
Classical stochastic approximation (SA) theory provides the analytical machinery—diffusion limits, Ornstein–Uhlenbeck linearization, Lyapunov stationary covariance, $\sqrt{t}$-CLT scaling—for analyzing such recursions once the gradient-noise covariance is specified. We do not claim novelty for that machinery. Our contribution is logically prior: we identify a structural property of the stochastic-gradient oracle that SA typically treats as exogenous, namely that the mini-batch gradient covariance is pinned down as $b^{-1}G^{\star}(\theta)$ under exchangeable sampling, where $G^{\star}(\theta)$ is the projected covariance of per-sample gradients. In likelihood settings this matrix reduces to Fisher information at the correctly specified point $\theta^{\star}$; for general losses it is the projected gradient covariance that, together with the Hessian $H^{\star}$, induces sandwich/Godambe geometry. Once $G^{\star}(\theta)$ is identified, the diffusion approximation for constant-step SGD is determined rather than modeled; the natural metric for rates is statistical rather than Euclidean; and batch size admits principled design rules under a fixed sampling budget. The subsequent Lyapunov analysis, rate bounds, and complexity guarantees follow by applying standard tools in this identified geometry.
@@ 1.1 Motivation: Simulation Optimization as Sampling Control
followed by $\theta^{+}=\theta-\eta\,g_{B}(\theta)$. Classical variance-reduction intuition focuses on the scalar variance of a component or of a scalarized performance measure, suggesting that larger $b$ should always help. In practice, under fixed sampling budgets, small $b$ often wins.
The key reason is geometric: the noise in $g_{B}(\theta)$ is not scalar. Its covariance is a matrix that encodes which directions are intrinsically informative and which are weakly identified by the underlying experiment. Our results show that, to leading order, this covariance scales as $b^{-1}G^{*}(\theta)$, where $G^{*}(\theta)$ is the projected covariance of per-sample gradients. For correctly specified likelihood objectives, $G^{*}(\theta^{\star})$ reduces to projected Fisher information at the well-specified point; for general losses $G^{*}(\theta)$ is the projected gradient covariance that, together with $H^{\star}$, induces sandwich/Godambe geometry. This provides the statistical foundation for interpreting SGD as a controlled diffusion whose effective temperature is $\tau=\eta/b$.
This sampling-control perspective is not limited to simulation optimization. Three canonical MOR problem classes share the same structure: (i) simulation optimization (queuing, inventory, revenue management), where each mini-batch gradient requires $b$ simulation replications and the replication budget is the binding constraint; (ii) stochastic programming and SAA, where scenario-based gradient estimates are drawn from an empirical distribution and the number of scenarios per iterate is a design variable subject to a total scenario budget; and (iii) online M-estimation in service systems, where per-transaction loss gradients arrive in a stream and the batch window trades off estimation quality against decision latency. In each case, sampling effort is explicit, $b$ is small relative to $d$, and the question “how many samples per iterate?” is operationally binding. The theory developed below applies uniformly to all three classes.
@@ 1.2 Core Insights
Before turning to formal statements, we summarize the main idea informally. When SGD computes a mini-batch gradient by averaging $b$ per-sample gradients, the resulting noise is not generic randomness. Its covariance—the matrix describing which directions are noisy and which are quiet—is determined by the loss function and the data-generating process.
This has three immediate consequences:
@@
    Mini-batch noise is anisotropic. Doubling the batch size reduces noise equally in all directions, but the shape of the noise ellipsoid is fixed by the identified covariance $G^{\star}$, with Fisher information recovered locally under correct specification.
@@
    The natural metric for local analysis is statistical, not Euclidean. Since the noise covariance is shaped by $G^{\star}$ near $\theta^{\star}$, measuring convergence in the corresponding statistical metric gives tighter and more interpretable bounds than the Euclidean metric.
@@
    Batch size is a design variable, not just a hyperparameter. Under a fixed oracle budget $N=Tb$, batch size controls the effective temperature $\tau=\eta/b$ of the induced diffusion. It governs the stationary covariance through the Lyapunov balance, making $b$ a design lever for effort allocation.
@@ 1.3 What We Show
(i) Identification from sampling theory.
@@
where $G^{*}(\theta)$ is the projected gradient covariance (Section 4). For correctly specified likelihood losses, $G^{*}(\theta^{\star})=F^{\star}(\theta^{\star})$.
@@
In the small-stepsize regime with $\tau=\eta/b$ held fixed, the SGD recursion admits a diffusion approximation whose volatility is determined by the identified covariance $G^{*}(\theta)$ (Section 5). Near a nondegenerate critical point, the limiting SDE linearizes to an OU process whose stationary covariance $\Sigma_{\infty}$ solves the Lyapunov equation
@@
making explicit that steady-state fluctuations are governed by the ratio of curvature to identified noise.
(iii) Matching rates and minimax lower bounds.
We analyze SGD performance in the frozen statistical geometry at $\theta^{\star}$. We establish $1/T$ upper bounds for mean-square error in the Fisher metric and match them with information-theoretic lower bounds via van Trees, showing that the rate and its dependence on intrinsic geometry are tight within the stated model class (Sections 6 and 7).
(iv) Oracle complexity in statistical conditioning.
The oracle complexity bounds scale with a statistical condition number and an effective dimension, isolating statistical difficulty from Euclidean conditioning. This clarifies when classical Hessian-based condition numbers can be misleading for statistical objectives.
@@
The framework above connects batch-size design to the identified statistical geometry; Section 9 discusses extensions to misspecification, dependence, and operational effort allocation.
@@
Theorem 4.3 — Identification of mini-batch covariance under exchangeable sampling: $\mathrm{Cov}(g_{B})=b^{-1}G^{\star}(\theta)+o(b^{-1})$. Theorem 5.4 — Diffusion approximation with volatility determined by $G^{\star}$. Thm. 5.6, Cor. 5.7 — OU linearization at $\theta^{\star}$; stationary covariance solves a Lyapunov equation. Thm. 6.5, Prop. 6.10 — Matching $1/(Tb)$ upper and lower bounds in the frozen Fisher metric. Theorem 7.3 — Oracle complexity $N\!=\!\widetilde{\Theta}(\kappa_{F}\,d_{\mathrm{eff}}/\varepsilon^{2})$ for Fisher-dual stationarity.
1.4 Organization
Section 2 reviews related work. Section 3 defines the oracle model and geometric objects. Section 4 derives the covariance identification from exchangeability. Section 5 builds the diffusion approximation, OU linearization, and discrete-time bridge. Section 6 proves matching upper and lower bounds in Fisher geometry, and Section 7 states oracle complexity. Section 8 provides numerical validation, and Section 9 discusses extensions and practical implications.
@@ 2 Literature Review and Positioning
We organize prior work along two axes: (i) whether the mini-batch noise covariance is assumed (exogenous) or derived from the sampling mechanism; and (ii) whether rates are stated in a Euclidean or in the identified statistical metric. Table 1 summarizes the resulting taxonomy.
Stochastic approximation and stationary covariance.
Classical SA theory studies recursions $\theta_{t+1}=\theta_{t}-\eta_{t}\{h(\theta_{t})+\xi_{t+1}\}$ and derives a $\sqrt{t}$-CLT whose asymptotic covariance is governed by a Lyapunov equation involving the Jacobian and a long-run noise covariance $\Gamma$ (Robbins and Monro, 1951; Dvoretzky, 1956; Kushner and Yin, 2003; Polyak and Juditsky, 1992). This framework takes $\Gamma$ as given; the OU linearization, Lyapunov balance, and $1/\sqrt{t}$ scaling are standard consequences. Our contribution is logically prior: we derive $\Gamma=G^{\star}(\theta)/b$ from the mini-batch sampling design (Theorem 4.3), converting the SA machinery from a conditional analysis into a structural one.
Diffusion approximations for SGD.
A substantial literature interprets constant-step SGD as a discretization of an SDE with temperature $\tau=\eta/b$ and various proposed diffusion matrices $\mathbf{C}(\theta)$ (Mandt et al., 2017; Li et al., 2017; Smith et al., 2021). In those works the diffusion covariance is either postulated, estimated from running iterates, or justified heuristically through moment matching. Our covariance identification removes this modeling degree of freedom: under exchangeable mini-batching, the diffusion matrix is determined as $G^{\star}(\theta)$ (Theorem 5.4), and the local OU regime inherits its Lyapunov balance at temperature $\tau$ (Corollary 5.7). Because the identification is logically prior to the discretization corrections of Li et al. (2017) or the Hessian-scaling regime of Smith et al. (2021), it is what enables the subsequent rate-matching bounds.
Scaling regimes.
Recent rigorous diffusion-limit work studies joint regimes where $b$ grows with $d$ (Paquette et al., 2024). Our analysis is complementary and closer to OR operating constraints: we keep $d$ fixed and treat $b$ as a design variable subject to a budget, with particular attention to the common regime $b\ll d$ where the matrix geometry of $G^{\star}(\theta)$ controls transient behavior and risk.
Information geometry, natural gradients, and Fisher-shaped preconditioning.
Information geometry treats Fisher information as a Riemannian metric and motivates natural-gradient updates $\theta_{t+1}=\theta_{t}-\eta\,F(\theta_{t})^{-1}\nabla L(\theta_{t})$ (Amari, 1998; Martens, 2020). Related methods—including stochastic gradient Fisher scoring (Ahn et al., 2012) and KFAC (Martens and Grosse, 2015)—inject or exploit Fisher geometry exogenously in the drift or preconditioner. Our contribution is distinct: even plain SGD, without any preconditioning, inherits the identified statistical geometry endogenously through its noise covariance under mini-batch sampling, changing the interpretation of diffusion strength, stationary risk, and batch-size tuning without requiring natural-gradient updates.
Minimax rates and information-theoretic limits.
Classical oracle lower bounds are stated in Euclidean norms under smoothness and strong convexity (Nemirovski et al., 2009; Agarwal et al., 2012), which may not reflect the intrinsic difficulty of statistical objectives. We state both upper and lower bounds in the identified statistical metric via a van Trees argument (Theorem 6.5 and Proposition 6.10), with complexity governed by statistical curvature and $d_{\operatorname{eff}}$ rather than ambient dimension.
OR context: sampling budgets.
In simulation optimization and stochastic programming, gradients are estimated from sampled scenarios and sampling cost is explicit (Fu, 2006; Shapiro et al., 2014; Chernoff, 1959). Because batching controls temperature $\tau=\eta/b$, it directly shapes diffusion amplitude and terminal risk; Section 9 outlines the resulting effort allocation problem.
Why small batches can work when $b\ll d$.
The empirical preference for small batches (Keskar et al., 2017; Shallue et al., 2019; McCandlish et al., 2018) is explained by our framework: smaller $b$ buys more contraction steps under fixed budgets, the identified covariance concentrates noise along statistically flat directions, and rates scale with $d_{\operatorname{eff}}\ll d$.
@@
Mini-batch noise covariance Exogenous / isotropic / heuristic $\Gamma$ Identified from sampling: $\mathrm{Cov}(g_{B})\propto b^{-1}G^{\star}(\theta)$; Fisher is a special case Sec. 4; Thm. 4.3
Diffusion matrix for SGD Postulated $\mathbf{C}(\theta)$ in SDEs Diffusion coefficient determined by sampling theory: $G^{\star}(\theta)$ Sec. 5; Thm. 5.4
OU / Lyapunov equilibrium Heuristic OU with ad hoc covariance OU linearization with Lyapunov balance driven by $G^{\star}$ at temperature $\tau=\eta/b$ Sec. 5; Cor. 5.7
Metric for rates Euclidean distance; $\kappa_{H}$ dominates Identified statistical metric is the natural risk Sec. 6; Thm. 6.5
Lower bounds Euclidean oracle lower bounds Information-theoretic lower bound in Fisher metric (van Trees) Sec. 6; Prop. 6.10